OpenStudy (anonymous):

Reviewing and need some refreshment help. Simplify: (8x^2+3x-9)-2(4x^2-8x+3)

sam (.sam.):

Any ideas?

OpenStudy (anonymous):

fix the second equation first?

sam (.sam.):

Yeah you distribute it what do you get?

OpenStudy (anonymous):

(8x^2-16x+6)

sam (.sam.):

Good, now use the first part subtract the second part (8x^2+3x-9)-(8x^2-16x+6)

OpenStudy (anonymous):

so next it would be (3x-9)-(-16x+6)?

sam (.sam.):

Nope, you subtract it like this \[(\color{red}{8x^2}\color{green}{+3x} \color{blue}{-9})-(\color{red}{8x^2} \color{green}{-16x} \color{blue}{+6})\] \[\color{red}{8x^2-8x^2=0} \\ \\ \color{green}{3x-(-16x)=19x} \\ \\ \color{blue}{-9-6=-15}\] You get \[19x-15\]

OpenStudy (anonymous):

Still got the same answer but i understand (: so would i find x to simplify or leave it like that?

sam (.sam.):

It says \(Simplify\) so you just leave it like this